OpenStudy (anonymous):

4x+12 over x^2-2x multiplied by x over 6x+18

OpenStudy (anonymous):

\[\frac{ 4x+12 }{ x ^{2}-2x }\times \frac{ x }{ 6x+18 }\]

OpenStudy (anonymous):

\[=\]...

OpenStudy (anonymous):

@Forever_Me @Jhannybean @RANE @kausarsalley

OpenStudy (anonymous):

Show all work, please and thank you ;3

OpenStudy (campbell_st):

do a little factoring \[\frac{4(x + 3)}{x(x -2)} \times \frac{x}{3(x + 3)}= \frac{ (x + 3) \times x \times 4}{(x + 3)\times x \times (x -2) \times 3}\] I've rearranged things so that you can see the common factors. whats left after removing the common factors in the answer.

OpenStudy (anonymous):

Could you simplfy it? Or show how?

OpenStudy (campbell_st):

ok whats common in the numerator and the denominator..?

OpenStudy (anonymous):

3

OpenStudy (campbell_st):

not quite you have (x + 3) and x that are in the top and bottom of the fraction. so what is left if you take them out of the fraction..?

OpenStudy (anonymous):

Oh, ok. 4?

OpenStudy (campbell_st):

you will have 4 in the top, now what about the bottom of the fraction take out (x + 3) and x... whats left...?

OpenStudy (anonymous):

(x-2)

OpenStudy (campbell_st):

and something else.. is left...

OpenStudy (anonymous):

x or the normal 3?

OpenStudy (campbell_st):

just the 3 so you are left with in compact form \[\frac{4x(x+ 3)}{3x(x+3)(x-2)} = \frac{4}{3(x - 2)}\] you could distribute the denominator for an answer as well

OpenStudy (anonymous):

Ok, like distribut the 4 over 3(x-2) o the other one?

OpenStudy (anonymous):

@campbell_st isn't it supposed to be 6(x+3) as your denominator (look at first step)

OpenStudy (anonymous):

??

OpenStudy (anonymous):

I got \[\frac{ 2 }{ 3(x-2) }\] for a final answer..

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

So it's not what @campbell_st came up with?

OpenStudy (anonymous):

i am sure because of the little mistake he made in the 1st step...

OpenStudy (anonymous):

......what he got is wrong

OpenStudy (campbell_st):

oops I made a mistake its should be \[\frac{4}{6(x -2)}\] which gives your answer... which is correct. just distribute the denominator \[\frac{2}{3x -6}\]

OpenStudy (anonymous):

Okay, but everything else is right?

OpenStudy (campbell_st):

well in the 1st factorised step you get \[\frac{4x(x + 3)}{6x(x+3)(x-2)}\] which gives your answer... then can be expanded as above.

OpenStudy (anonymous):

is this right for number 11?

OpenStudy (campbell_st):

yep its a solution... you may need the factorised step above... but it depends on what your teacher wants.

OpenStudy (anonymous):

@campbell_st please give @kausarsalley a medal

OpenStudy (anonymous):

ok where should i pput that astep?

OpenStudy (anonymous):

Where do i put that step @campbell_st ?